Lance Armstrong (fully doped) had a maximum oxygen consumption (VO2) of 5.5 liters O2/min. If during maximum exertion his arterial oxygen content is 20 ml/100 ml blood and his venous content is 2 ml/100ml blood, a) calculate his cardiac output. In mammals in general, blood volume is equal to 7% of body mass (assume one liter of blood weighs 1.0 kg). If he weighs 65kg b) during this maximum exertion how long will it take to circulate his entire blood volume (i.e., how long for heart to pump his blood volume once?)

Potentially useful information:

PV = nRT, one mole of STP gas occupies 22.4L. Saturation water vapor at 37°C is 47 mmHg, at 25°C is 24 mmHg and at -30°C is 0.4 mmHg. Atmospheric air is 21% O2. Sea level barometric pressure (Pb) is 760 mmHg.

PAO2 = 0.21 (Pb - 47) – PaCO2/RER (A= arterial a =alveolar)

Note RER = respiratory exchange ratio = CO2 produced / O2 consumed

normal value: PaCO2 = 40 mmHg

Fick equation: VO2= CO(CaO2 - CvO2) where CO = cardiac output.

**Use textbook RER values for fat: 0.7, Protein: 0.8 and Carbs: 1
Answer #1

I HOPE ITS HELP FUL TO YOU..IF YOU HAVE ANY DOUBTS PLS COMMENTS BELOW..I WILL BE THERE TO HELP YOU..ALL THE BEST...PLS RATE THUMBS UP....

AS FOR GIVEN DATA...

Lance Armstrong (fully doped) had a maximum oxygen consumption (VO2) of 5.5 liters O2/min. If during maximum exertion his arterial oxygen content is 20 ml/100 ml blood and his venous content is 2 ml/100ml blood, a) calculate his cardiac output. In mammals in general, blood volume is equal to 7% of body mass (assume one liter of blood weighs 1.0 kg). If he weighs 65kg b) during this maximum exertion how long will it take to circulate his entire blood volume (i.e., how long for heart to pump his blood volume once?)

Potentially useful information:

PV = nRT, one mole of STP gas occupies 22.4L. Saturation water vapor at 37°C is 47 mmHg, at 25°C is 24 mmHg and at -30°C is 0.4 mmHg. Atmospheric air is 21% O2. Sea level barometric pressure (Pb) is 760 mmHg.

PAO2 = 0.21 (Pb - 47) – PaCO2/RER (A= arterial a =alveolar)

Note RER = respiratory exchange ratio = CO2 produced / O2 consumed

normal value: PaCO2 = 40 mmHg

Fick equation: VO2= CO(CaO2 - CvO2) where CO = cardiac output.

**Use textbook RER values for fat: 0.7, Protein: 0.8 and Carbs: 1

ANS ::-

EXPLANATION ::-

(1)

Cardiac Output can be calculated using Fick equation.

Fick equation:VO2 = CO(CaO2 - CvO2)

CO : cardiac output

VO2: mximun oxygen taken up

CaO2 : Arterial oxygen content or oxygen concentration of arterial blood

CvO2 : Venous oxygen content or Oxygen concen tration at mixed venous blood

(CaO2 - CvO2) : Arterial venous oxygen difference

CO = V O2/ (CaO2 - CvO2)

= 5.5 x 10^3 ml O2/min / ( 20 ml/100 ml blood) - ( 2 ml/100ml blood) = 5.5 x 10^3 ml O2/min / 18 ml/100 ml blood = 0.305 x 10^3 ml blood/min.

1 L = 1000ml

5.5 L = 5.5 x 10^3 ml

(2)

From Fick equation we know that,

VO2 = CO (CaO2 - CaO2) where CO = Cardiac output ; VO2 = Volume of oxygen consumed per minute= 5.5 L/min ; CaO2 = concentration of oxygen in arterial blood = 20mL/100mL = 200mL/ 1L ;  CaO2 = concentration of oxygen in veneous blood = 2ml/100ml = 20mL/1L

Putting these values in the above equatio we get, CO = 5500ml/(200-20) = 30.55 L/min = cardiac output

Blood volume = 0.07 x 65 = 4.55kg = 4.55 L of blood.

b) During maximum exertion the time taken to circulate his entire blood volume is, (Blood Volume)/(CO) = 4.55/30.55 = 0.148 min = 8.88 seconds

I HOPE YOU UNDERSTAND ...I HOPE ITS HELP FUL TO YOU..PLS RATE THUMBS UP ITS HELPS ME ALOT ...

THANK YOU....!!